# Upper Approximation Bounds for Neural Oscillators

### v1

@@ Abstract
Neural oscillators, originating from the second-order ordinary differential equations (ODEs), have demonstrated competitive performance in stably learning causal mappings between long-term sequences or continuous temporal functions. However, theoretically quantifying the capacities of their neural network architectures remains a significant challenge. In this study, the neural oscillator consisting of a second-order ODE followed by a multilayer perceptron (MLP) is considered. Its upper approximation bound for approximating causal and uniformly continuous operators between continuous temporal function spaces and that for approximating uniformly asymptotically incrementally stable second-order dynamical systems are derived. The established proof method of the approximation bound for approximating the causal continuous operators can also be directly applied to state-space models consisting of a linear time-continuous complex recurrent neural network followed by an MLP. Theoretical results reveal that the approximation error of the neural oscillator for approximating the second-order dynamical systems scales polynomially with the reciprocals of the widths of two utilized MLPs, thus mitigating the curse of parametric complexity. The decay rates of two established approximation error bounds are validated through two numerical cases. These results provide a robust theoretical foundation for the effective application of the neural oscillator in science and engineering.
@@
journal: Neural Networks
@@ 1 Introduction
Modeling the mapping relationships of long sequences or continuous temporal functions is a significant challenge in machine learning. A large variety of neural network architectures have been proposed to address sequential and functional mapping tasks, including recurrent neural networks (RNNs) (Rumelhart et al., 1986) and their variants (Hochreiter and Schmidhuber, 1997; Cho et al., 2014), attention-based architectures (Vaswani et al., 2017), state-space (SS) models based on the first-order ordinary differential equations (ODEs) (Gu et al., 2020, 2021, 2022), and neural oscillators based on the second-order ODEs (Rusch and Mishra, 2020, 2021; Lanthaler et al., 2023; Rusch and Rus, 2024; Huang and Xia, 2025). SS models and neural oscillators possess the fast inference capability of RNNs together with the competitive performance of the attention-based architectures in long-range sequence learning (Gu and Dao, 2023; Rusch and Rus, 2024). Furthermore, neural oscillators have rigorously demonstrated to mitigate the problem of vanishing or exploding gradients (Rusch and Mishra, 2020, 2021) and can provide stable dynamics in learning long-range interactions (Rusch and Rus, 2024). A recent overview on SS models and neural oscillators was provided by Tiezzi et al. (2025).
@@
Despite the practical success of SS models and neural oscillators, advancing their theoretical understanding, particularly in quantifying the capacities of their network architectures, remains a significant challenge. For the universal approximation properties of SS models, Li et al. (2022) demonstrated that linear RNNs can universally approximate linear, causal, continuous, regular, and time-homogeneous operators between continuous temporal function spaces along with an approximation bound. Based on the result of Li et al. (2022) and the Volterra series (Boyd et al., 1984), Wang and Xue (2023) proved that given sufficiently large model sizes, an SS model consisting of a time-continuous linear RNN followed by a polynomial nonlinear projection can approximate any time-homogeneous causal continuous operator between continuous temporal function spaces. Similarly, Muca Cirone et al. (2024) demonstrated that a selective SS model comprising a linear controlled differential equation followed by a multilayer perceptron (MLP) can universally approximate causal continuous operators of continuous temporal functions. However, neither result provides a quantitative relationship between the approximation error and the number of parameters required in the SS models. Orvieto et al. (2024) derived an upper approximation bound for an SS model consisting of a time-discrete linear RNN followed by an MLP. However, this result is only limited to the mapping relationships between finite-length sequences and does not apply to infinite-length sequences or continuous temporal functions. In addition to above results, the approximation property of the original RNN for approximating the first-order dynamical systems (Hanson and Raginsky, 2020) and those for approximating fading memory causal operators of time-discrete sequences (Grigoryeva and Ortega, 2018a, b; Gonon et al., 2023) have been studied.
@@
For neural oscillators, a fundamental lemma proposed by Lanthaler et al. (2023) provides a novel insight that the time-varying sine transform coefficients of input continuous temporal functions can be used to construct a universal neural network architecture for the causal continuous operators between continuous temporal function spaces. Based on this fundamental lemma, the neural oscillator architecture consisting of a second-order ODE followed by an MLP was proven to be a universal approximator for such causal continuous operators (Lanthaler et al., 2023; Rusch and Rus, 2024). However, these theoretical results are only qualitative, and the quantitative approximation bounds of neural oscillators remain undeveloped.
@@
In this study, the neural oscillator consisting of a second-order ODE followed by an MLP is considered. Its upper approximation bound for approximating causal and uniformly continuous operators between continuous temporal function spaces and that for approximating uniformly asymptotically incrementally stable second-order dynamical systems are derived. In Section 2, the network architecture of the neural oscillator is briefly introduced, and the newly established theoretical approximation bounds of the neural oscillator are presented informally to facilitate understanding. The formal statements of two theorems that establish the approximation bounds for the neural oscillator, accompanied by the supporting lemmas essential for their proofs, are provided in Section 3. In Section 4, the decay rates of the two established approximation error bounds are validated through two numerical cases. The main contributions of this study are summarized as follows:
@@
    During the mathematical derivation in Section 3.2, Lemma 4 enhances the proof of a fundamental Lemma 3.1.2 on differential equation flow by Van Handel (2007). The results in Lemma 6 and D demonstrate that two methods, one based on the differential equation flow and the other one on the Grönwall’s inequality (Ames and Pachpatte, 1997), are equivalent in establishing the bound of the difference between the solutions of two second-order ODEs driven by the same input functions. The proof of Theorem 2 establishes a new method to ensure that all possible solutions of a neural ODE, whose MLP is employed to approximate the derivative function of a target ODE, remain within a domain where the error between the MLP of the neural ODE and the derivative function of the target ODE can be quantified.
@@
Notations: In this study, the $\it{L^{r}}$-norm of an n-dimensional vector $\mathbf{a}=[a_{1},a_{2},...,a_{n}]^{\top},1\leq r\leq+\infty$, is $|\mathbf{a}|_{r}=[\sum_{i=1}^{n}|a_{i}|^{r}]^{1/r}$ for $r<\infty$, or $|\mathbf{a}|_{r}=\max_{1\leq i\leq n}|a_{i}|$ for $r=+\infty$. $|\mathbf{a}|$ is short for $|\mathbf{a}|_{1}$. For a matrix $\mathbf{M}=[m_{ij}]_{p\times q}$, its $L^{r,s}$-norm, $1\leq r,s\leq+\infty$, is $\lvert\mathbf{M}\rvert_{r,s}=\lvert\left[\lvert\mathbf{m}_{1}\rvert_{r},\lvert\mathbf{m}_{2}\rvert_{r},\dots,\lvert\mathbf{m}_{q}\rvert_{r}\right]^{\top}\rvert_{s}$, where $\mathbf{m}_{j}$ is the $j^{\text{th}}$ column vector of the matrix $\mathbf{M}$. The $L^{r}$-norm of $\mathbf{M}$ is $|\mathbf{M}|_{r}=\sup_{|\mathbf{a}|_{r}=1}|\mathbf{Ma}|_{r}/|\mathbf{a}|_{r}$. Re($\cdot$) and Im($\cdot$) represent the real and imaginary parts of a complex number, respectively. $\lfloor\cdot\rfloor$ and $\lceil\cdot\rceil$ represent the floor and ceiling functions, respectively. $\mathbb{Z}$ represents the integer space. $\mathbb{R}$ represents the real number space. $\mathbb{R}^{p{\times}q}$ represents the $p\times q$-dimensional real matrix space. $f^{\prime}(t)=\mathrm{d}f(t)/\mathrm{d}t$, $f^{\prime\prime}(t)=\mathrm{d}^{2}f(t)/\mathrm{d}t^{2}$, and $f^{\prime\prime\prime}(t)=\mathrm{d}^{3}f(t)/\mathrm{d}t^{3}$. $C_{0}([0,T];\mathbb{R}^{p})$ represents a p-dimensional vector-valued real continuous function space defined over $[0,T]$ with respect to $L^{\infty}$-norm ${\lVert\cdot\rVert}_{L^{\infty}}$, e.g., ${\lVert\mathbf{u}(t)\rVert}_{L^{\infty}}$ $\sup_{t\in[0,T]}|\mathbf{u}(t)|$ for $\mathbf{u}(t)=[u_{1}(t),u_{2}(t),...,u_{p}(t)]^{\top}\in C_{0}([0,T];{\mathbb{R}^{p}})$. In addition, $L^{\infty}{[0,t_{0}]}$-norm ${\lVert\cdot\rVert}_{L^{\infty}[0,t_{0}]}$ is ${\lVert\mathbf{u}(t)\rVert}_{L^{\infty}[0,t_{0}]}=\sup_{t\in[0,t_{0}]}|\mathbf{u}(t)|$ and $L^{2}$-norm ${\lVert\cdot\rVert}_{L^{2}}$ is ${\lVert\mathbf{u}(t)\rVert}_{L^{2}}=\sqrt{\int_{0}^{T}|\mathbf{u}(t)|_{2}^{2}\mathrm{d}t}$. Big $O$ notation: $f(n)=O\left[g(n)\right]$ if there exist positive constants $c$ and $n_{0}$ such that $f(n)\leq c\cdot g(n)$ for all $n\geq n_{0}$.
@@ 2 Neural Oscillator
Suppose $\mathbf{u}(t)=[u_{1}(t),u_{2}(t),...,u_{p}(t)]^{\top}\in C_{0}([0,T],\mathbb{R}^{p})$, a neural oscillator architecture mapping $\mathbf{u}(t)$ to $\mathbf{y}(t)=[y_{1}(t),y_{2}(t),...,y_{q}(t)]^{\top}$ over $t\in[0,T]$ is
@@
$\mathbf{x}(t)=[x_{1}(t),x_{2}(t),...,x_{r}(t)]^{\top}$ is an intermediate r-dimensional real-valued state function, $\mathbf{W}_{i}$, i = 1 and 2, $\tilde{\mathbf{W}}_{j}$, j = 1, 2,…, $h_{\it{\Pi}}$, are real-valued tunable weight matrices, $\mathbf{b}_{i}$ and $\tilde{\mathbf{b}}_{j}$ are real-valued tunable bias vectors, $h_{\it{\Pi}}$ is the depth of ${\it{\Pi}}(\cdot)$, and $\sigma_{\it{\Gamma}}\left(\cdot\right)$ and $\sigma_{\it{\Pi}}\left(\cdot\right)$ are the activation functions of ${\it{\Gamma}}(\cdot)$ and ${\it{\Pi}}(\cdot)$, respectively. The neural oscillator architecture in Eq. (1) can be denoted as $\mathbf{y}(t)={\it{\Pi}}\circ{\Phi}_{\it{\Gamma}}[\mathbf{u}(\tau)](t)={\it{\Pi}}\left\{{{\Phi}_{\it{\Gamma}}[{\mathbf{u}(\tau)}](t),\mathbf{u}(0),t}\right\}$, where $\circ$ is the composition operator and $\mathbf{x}(t)=\Phi_{\it{\Gamma}}\left[\mathbf{u}(\tau)\right](t)$ represents the casual operator characterized by the second-order ODE in Eq. (1).
When $\sigma_{\it{\Gamma}}(\cdot)$ is the Rectified Linear Unit (ReLU) $\sigma_{\mathrm{ReLU}}(x)=\mathrm{max}(0,x)$ (Nair and Hinton, 2010), Eq. (1) can function as a universal approximator of causal continuous operators by first encoding an input function $\mathbf{u}(t)$ into its time-varying sine transform coefficients through a linear second-order ODE in Eq. (1), and subsequently mapping these coefficients, together with the initial value $\mathbf{u}(0)$ of $\mathbf{u}(t)$ and time t, to an output function, as established in Theorem 1 of this study. This capability is similar to the universal approximation property of neural operators (Kovachki et al., 2023). Moreover, for the second-order dynamical systems in science and engineering (Strogatz, 2024), Eq. (1) can directly approximate these systems using its general ODE with an efficient neural network size, as demonstrated in Theorem 2 of this study.
@@
The initial value u(0) added in ${\it{\Pi}}(\cdot)$ eliminates the need for the assumption of $\mathbf{u}(0)=\mathbf{0}$ in the previous universal approximation theorems for neural oscillators, as presented in Lanthaler et al. (2023) and Rusch and Rus (2024). The influence of time in non-autonomous causal operators is characterized by a term $0.25t^{2}$ in Lanthaler et al. (2023), which arises as the solution to a second-order ODE $f^{\prime\prime}(t)=0.5$. However, this has the possibility of causing unbounded derivatives of the MLP ${\it{\Pi}}(\cdot)$. For example, if ${\it{\Pi}}(t^{2})=g(t)$ and $g^{\prime}(t)$ is continuous and bounded over $t\in[0,T]$ with $g^{\prime}(0)\neq 0$, then $g^{\prime}(t)=2t{\it{\Pi}}^{\prime}(t^{2})$ and ${\it{\Pi}}^{\prime}(x)$ tends to infinity as x approaches zero. The potential unbounded derivatives of ${\it{\Pi}}(\cdot)$ may induce challenges for subsequent theoretical analysis. To address this problem, in this study, a term t is directly added to ${\it{\Pi}}(\cdot)$ to ensure that the derivative of ${\it{\Pi}}(\cdot)$ with respect to t can be bounded.
@@
Assumption 1: $C_{0}([0,T];{\mathbb{R}}^{p})$ with $L^{\infty}$-norm is a Banach space. Its metric is defined as ${\lVert{\mathbf{u}_{1}(t)-\mathbf{u}_{2}(t)}\rVert}_{L^{\infty}}$ with $\mathbf{u}_{1}(t)$ and $\mathbf{u}_{2}(t)$ $\in$ $C_{0}([0,T];\mathbb{R}^{p})$.
Assumption 2: K is a compact subset of $C_{0}([0,T];\mathbb{R}^{p})$} and it is closed. By virtue of the Arzelà-Ascoli theorem (Beattie and Butzmann, 2013), all functions belonging to K are uniformly bounded. There exists a constant $B_{K}$ such that all elements of $K$ are bounded by $B_{K}$, that is ${\lVert{u_{i}(t)}\rVert}_{L^{\infty}}\leq B_{K}$ for all $\mathbf{u}(t)\in K$ and $1\leq i\leq p$, where $u_{i}(t)$ is the $i^{\mathrm{th}}$ element of $\mathbf{u}(t)$. It follows from the Arzelà-Ascoli theorem that the functions belonging to $K$ are equicontinuous, and there exists a continuous modulus of continuity $\mathbf{\upphi}_{K}(t)=[\phi_{K,1}(t),\phi_{K,2}(t),...,\phi_{K,p}(t)]^{\top}$ such that for arbitrary $0\leq t_{1}\leq t_{2}\leq T$ and all $\mathbf{u}(t)\in K$, it satisfies $|u_{i}(t_{1})-u_{i}(t_{2})|\leq\phi_{K,i}(t_{2}-t_{1})$, thereby $|\mathbf{u}(t_{1})-\mathbf{u}(t_{2})|\leq\left|\mathbf{\upphi}_{K}(t_{2}-t_{1})\right|$, where $\phi_{K,i}(t):[0,+\infty)\to[0,+\infty)$ is monotonic with $\phi_{K,i}(t)\to 0$ as $t\to 0$.

### v2

@@ Abstract
Neural oscillators, originating from second-order ordinary differential equations (ODEs), have demonstrated strong performance in stably learning causal mappings between long-term sequences or continuous temporal functions, as well as in accurately approximating physical systems. However, theoretically quantifying the capacities of their neural network architectures remains a significant challenge. In this study, the neural oscillator consisting of a second-order ODE followed by a multilayer perceptron (MLP) is considered. Its upper approximation bound for approximating causal and uniformly continuous operators between continuous temporal function spaces and that for approximating uniformly asymptotically incrementally stable second-order dynamical systems are derived. The established proof method of the approximation bound for approximating the causal continuous operators can also be directly applied to state-space models consisting of a linear time-continuous complex recurrent neural network followed by an MLP. Theoretical results reveal that the approximation error of the neural oscillator for approximating the second-order dynamical systems scales polynomially with the reciprocals of the widths of two utilized MLPs, thus overcoming the curse of parametric complexity. The convergence rates of two established approximation error bounds are validated through four numerical cases. These results provide a robust theoretical foundation for the effective application of the neural oscillator in science and engineering.
@@
journal: Mechanical Systems and Signal Processing
@@ 1 Introduction
Modeling the mapping relationships of long sequences or continuous temporal functions is a significant challenge in machine learning. A large variety of neural network architectures have been proposed to address sequential and functional mapping tasks, including recurrent neural networks (RNNs) (Rumelhart et al., 1986) and their variants (Hochreiter and Schmidhuber, 1997; Cho et al., 2014), attention-based architectures (Vaswani et al., 2017), state-space (SS) models based on the first-order ordinary differential equations (ODEs) (Gu et al., 2020, 2021, 2022), and neural oscillators based on the second-order ODEs (Rusch and Mishra, 2020, 2021; Lanthaler et al., 2023; Rusch and Rus, 2024; Huang and Xia, 2025). SS models and neural oscillators possess the fast inference capability of RNNs together with the competitive performance of the attention-based architectures in long-range sequence learning (Gu and Dao, 2023; Rusch and Rus, 2024). Furthermore, neural oscillators have rigorously demonstrated to mitigate the problem of vanishing or exploding gradients (Rusch and Mishra, 2020, 2021). They can provide stable dynamics in learning long-range interactions (Rusch and Rus, 2024).
Regarding the applications of neural oscillators, Rusch and Rus (2024) utilized two neural oscillator models to predict heart rates using real-world data from wrist-worn devices, and showed that the neural oscillators outperformed some SS models. Stölzle and Della Santina (2024) employed an input-to-state stable neural oscillator to characterize the latent dynamic equation of an autoencoder framework, and the proposed model was utilized to learn complex nonlinear dynamics of a continuum soft robot directly from raw pixel images. An integral-saturated proportional–integral–derivative controller with potential-energy-shaping feedforward compensation was developed to generate corrective actuation commands that drive the robot toward desired trajectories, using raw pixel images as the sole sensing modality. Kapoor et al. (2024) employed the neural oscillators to enhance the generalization of physics-informed machine learning beyond the training domain. By exploiting the inherent causality and temporal sequential characteristics of partial differential equation (PDE) solutions, this work fused a physics-informed machine learning model with the neural oscillators to process the model outputs as sequential data and predict solutions in unseen temporal regions. The effectiveness of this framework was demonstrated across multiple time-dependent nonlinear PDEs. Huang and Xia (2025) numerically demonstrated that neural oscillators can directly learn the nonlinear derivative terms governing complex ODE systems. A recent overview on SS models and neural oscillators was provided by Tiezzi et al. (2025).
@@
Despite the practical success of SS models and neural oscillators, advancing their theoretical understanding, particularly in quantifying the capacities of their network architectures, remains a significant challenge. For the universal approximation properties of SS models, Li et al. (2022) demonstrated that linear RNNs can universally approximate linear, causal, continuous, regular, and time-homogeneous operators between continuous temporal function spaces along with an approximation bound. Based on the result of Li et al. (2022) and the Volterra series (Boyd et al., 1984), Wang and Xue (2023) proved that given sufficiently large model sizes, an SS model consisting of a time-continuous linear RNN followed by a polynomial nonlinear projection can approximate any time-homogeneous causal continuous operator between continuous temporal function spaces. Similarly, Muca Cirone et al. (2024) demonstrated that a selective SS model comprising a linear controlled differential equation followed by a multilayer perceptron (MLP) can universally approximate causal continuous operators of continuous temporal functions. However, neither result provides a quantitative relationship between the approximation error and the number of parameters required in the SS models. Orvieto et al. (2024) derived an upper approximation bound for an SS model consisting of a discrete-time linear RNN followed by an MLP. However, this result is only limited to the mapping relationships between finite-length sequences and does not apply to infinite-length sequences or continuous temporal functions. In addition to above results, the approximation property of the original RNN for approximating the first-order dynamical systems (Hanson and Raginsky, 2020) and that for approximating fading memory causal operators of discrete-time sequences (Grigoryeva and Ortega, 2018a, b; Gonon et al., 2023) have been studied.
@@
For the neural oscillators, a fundamental lemma proposed by Lanthaler et al. (2023) provides a novel insight that the time-varying sine transform coefficients of time-continuous input functions can be used to construct a universal neural network architecture for the causal continuous operators between continuous temporal function spaces. Based on this fundamental lemma, the neural oscillator architecture consisting of a second-order ODE followed by an MLP was proven to be a universal approximator for such causal continuous operators (Lanthaler et al., 2023; Rusch and Rus, 2024). However, these theoretical results are only qualitative, and the quantitative approximation bounds of the neural oscillators remain undeveloped.
@@
In this study, the neural oscillator consisting of a second-order ODE followed by an MLP is considered. Its upper approximation bound for approximating causal and uniformly continuous operators between continuous temporal function spaces and that for approximating uniformly asymptotically incrementally stable second-order dynamical systems are derived. In Section 2, the network architecture of the neural oscillator is briefly introduced, and the newly established theoretical approximation bounds of the neural oscillator are presented informally to facilitate understanding. The formal statements of two theorems that establish the approximation bounds for the neural oscillator, accompanied by the supporting lemmas essential for their proofs, are provided in Section 3. In Section 4, the convergence rates of the two established approximation error bounds are validated through four numerical cases. The main contributions of this study are summarized as follows:
@@
    During the mathematical derivation in Section 3.2, Lemma 4 enhances the proof of a fundamental Lemma 3.1.2 on differential equation flow by Van Handel (2007). The results in Lemma 6 and D demonstrate that two methods, one based on the differential equation flow and the other one on the Grönwall’s inequality (Ames and Pachpatte, 1997), are equivalent in establishing the bound of the difference between the solutions of two second-order ODEs driven by the same input functions. The proof of Theorem 2 establishes a new method to ensure that all possible solutions of a neural ODE, whose MLP is employed to approximate the derivative term function of a target ODE, remain within a domain where the error between the MLP of the neural ODE and the derivative term function of the target ODE can be quantified.
@@
Notations: In this study, the $\it{L^{r}}$-norm of an n-dimensional vector $\mathbf{a}=[a_{1},a_{2},...,a_{n}]^{\top},1\leq r\leq+\infty$, is $|\mathbf{a}|_{r}=[\sum_{i=1}^{n}|a_{i}|^{r}]^{1/r}$ for $r<\infty$, or $|\mathbf{a}|_{r}=\max_{1\leq i\leq n}|a_{i}|$ for $r=+\infty$. $|\mathbf{a}|$ is short for $|\mathbf{a}|_{1}$. For a matrix $\mathbf{M}=[m_{ij}]_{p\times q}$, its $L^{r,s}$-norm, $1\leq r,s\leq+\infty$, is $\lvert\mathbf{M}\rvert_{r,s}=\lvert\left[\lvert\mathbf{m}_{1}\rvert_{r},\lvert\mathbf{m}_{2}\rvert_{r},\dots,\lvert\mathbf{m}_{q}\rvert_{r}\right]^{\top}\rvert_{s}$, where $\mathbf{m}_{j}$ is the $j^{\text{th}}$ column vector of $\mathbf{M}$. The $L^{r}$-norm of $\mathbf{M}$ is $|\mathbf{M}|_{r}=\sup_{|\mathbf{a}|_{r}=1}|\mathbf{Ma}|_{r}/|\mathbf{a}|_{r}$. Re($\cdot$) and Im($\cdot$) represent the real and imaginary parts of a complex number, respectively. $\lfloor\cdot\rfloor$ and $\lceil\cdot\rceil$ represent the floor and ceiling functions, respectively. $\mathbb{Z}$ represents the integer space. $\mathbb{R}$ represents the real number space. $\mathbb{R}^{p{\times}q}$ represents the $p\times q$-dimensional real matrix space. $f^{\prime}(t)=\mathrm{d}f(t)/\mathrm{d}t$, $f^{\prime\prime}(t)=\mathrm{d}^{2}f(t)/\mathrm{d}t^{2}$, and $f^{\prime\prime\prime}(t)=\mathrm{d}^{3}f(t)/\mathrm{d}t^{3}$. $C_{0}([0,T];\mathbb{R}^{p})$ represents a p-dimensional vector-valued real continuous function space defined over $[0,T]$ under the $L^{\infty}$-norm ${\lVert\cdot\rVert}_{L^{\infty}}$, e.g., ${\lVert\mathbf{u}(t)\rVert}_{L^{\infty}}$ $\sup_{t\in[0,T]}|\mathbf{u}(t)|$ for $\mathbf{u}(t)=[u_{1}(t),u_{2}(t),...,u_{p}(t)]^{\top}\in C_{0}([0,T];{\mathbb{R}^{p}})$. In addition, the $L^{\infty}{[0,t_{0}]}$-norm ${\lVert\cdot\rVert}_{L^{\infty}[0,t_{0}]}$ is ${\lVert\mathbf{u}(t)\rVert}_{L^{\infty}[0,t_{0}]}=\sup_{t\in[0,t_{0}]}|\mathbf{u}(t)|$ and the $L^{2}$-norm ${\lVert\cdot\rVert}_{L^{2}}$ is ${\lVert\mathbf{u}(t)\rVert}_{L^{2}}=\sqrt{\int_{0}^{T}|\mathbf{u}(t)|_{2}^{2}\mathrm{d}t}$. $f(n)=O\left[g(n)\right]$ if there exist positive constants $c$ and $n_{0}$ such that $f(n)\leq c\cdot g(n)$ for all $n\geq n_{0}$. $f(n)={\it{\Omega}}\left[g(n)\right]$ if there exist positive constants $c$ and $n_{0}$ such that $f(n)\geq c\cdot g(n)$ for all $n\geq n_{0}$.
@@ 2 Neural Oscillator
For a given $\mathbf{u}(t)=[u_{1}(t),u_{2}(t),...,u_{p}(t)]^{\top}\in C_{0}([0,T],\mathbb{R}^{p})$, a neural oscillator architecture mapping $\mathbf{u}(t)$ to $\mathbf{y}(t)=[y_{1}(t),y_{2}(t),...,y_{q}(t)]^{\top}$ over $t\in[0,T]$ is
@@
$\mathbf{x}(t)=[x_{1}(t),x_{2}(t),...,x_{r}(t)]^{\top}$ is an intermediate r-dimensional real-valued state function, $\mathbf{W}_{i}$, i = 1 and 2, $\tilde{\mathbf{W}}_{j}$, j = 1, 2,…, $h_{\it{\Pi}}$, are real-valued trainable weight matrices, $\mathbf{b}_{i}$ and $\tilde{\mathbf{b}}_{j}$ are real-valued trainable bias vectors, $h_{\it{\Pi}}$ is the depth of ${\it{\Pi}}(\cdot)$, and $\sigma_{\it{\Gamma}}\left(\cdot\right)$ and $\sigma_{\it{\Pi}}\left(\cdot\right)$ are the activation functions of ${\it{\Gamma}}(\cdot)$ and ${\it{\Pi}}(\cdot)$, respectively. The neural oscillator architecture in Eq. (1) can be denoted as $\mathbf{y}(t)={\it{\Pi}}\circ{\Phi}_{\it{\Gamma}}[\mathbf{u}(\tau)](t)={\it{\Pi}}\left\{{{\Phi}_{\it{\Gamma}}[{\mathbf{u}(\tau)}](t),\mathbf{u}(0),t}\right\}$, where $\circ$ is the composition operator and $\mathbf{x}(t)=\Phi_{\it{\Gamma}}\left[\mathbf{u}(\tau)\right](t)$ represents the casual operator characterized by the second-order ODE in Eq. (1).
@@
The initial value u(0) added in ${\it{\Pi}}(\cdot)$ eliminates the need for the assumption of $\mathbf{u}(0)=\mathbf{0}$ in the previous universal approximation theorems for the neural oscillators, as presented in Lanthaler et al. (2023) and Rusch and Rus (2024). The influence of time in non-autonomous causal operators is characterized by a term $0.25t^{2}$ in Lanthaler et al. (2023), which arises as the solution to a second-order ODE $f^{\prime\prime}(t)=0.5$. However, this has the possibility of causing unbounded derivatives of the MLP ${\it{\Pi}}(\cdot)$. For example, if ${\it{\Pi}}(t^{2})=g(t)$ and $g^{\prime}(t)$ is continuous and bounded over $t\in[0,T]$ with $g^{\prime}(0)\neq 0$, then $g^{\prime}(t)=2t{\it{\Pi}}^{\prime}(t^{2})$ and ${\it{\Pi}}^{\prime}(x)$ tends to infinity as x approaches zero. The potential unbounded derivatives of ${\it{\Pi}}(\cdot)$ may induce challenges for subsequent theoretical analysis. To address this problem, in this study, a term t is directly added to ${\it{\Pi}}(\cdot)$ to ensure that the derivative of ${\it{\Pi}}(\cdot)$ with respect to t can be bounded.
Figure 1: The two-module structure and error propagation pathways of the neural oscillator.
As illustrated in Figure 1, the neural oscillator is a two-module structure. The first module is a second-order neural ODE encoder characterized by ${\it{\Gamma}}(\cdot)$ that encodes the input function $\mathbf{u}(t)$ into the intermediate state function $\mathbf{x}(t)$. The second module is a readout MLP ${\it{\Pi}}(\cdot)$ that maps $\mathbf{x}(t)$, together with the initial value $\mathbf{u}(0)$ of $\mathbf{u}(t)$ and time t, to the output function $\mathbf{y}(t)$. As established in Theorem 1 of this study, when $\sigma_{\it{\Gamma}}(\cdot)$ is the Rectified Linear Unit (ReLU) $\sigma_{\mathrm{ReLU}}(x)=\mathrm{max}(0,x)$ (Nair and Hinton, 2010), the neural oscillator can function as a universal approximator of causal continuous operators. In Theorem 1, the total approximation error is decomposed into two sequential errors. The first arises from truncating the sine transform coefficients of $\mathbf{u}(t)$, and the second stems from approximating the target mapping from $\mathbf{u}(0)$, t, and the sine transform coefficients of $\mathbf{u}(t)$ to $\mathbf{y}(t)$ by ${\it{\Pi}}(\cdot)$. Moreover, for the second-order dynamical systems in science and engineering (Strogatz, 2024), the neural oscillator can leverage its MLPs ${\it{\Gamma}}(\cdot)$ and ${\it{\Pi}}(\cdot)$ to directly approximate the derivative term and output functions of these systems with efficient neural network sizes, as demonstrated in Theorem 2 of this study. In this situation, the total approximation error again consists of two sequential errors. The first is induced by approximating a target second-order ODE by the neural ODE characterized by ${\it{\Gamma}}(\cdot)$, and the second is caused by approximating a target output function by ${\it{\Pi}}(\cdot)$.
Traditional neural ODE models, including the first-order and second-order neural ODEs (Chen et al., 2018; Norcliffe et al., 2020), leverage the relationships between their initial conditions and solutions at specific time instants to universally approximate finite-dimensional continuous functions (Celledoni et al., 2023). The neural oscillator augments a second-order neural ODE with a readout MLP. It acts as a universal approximator for causal continuous operators by first encoding input functions into a finite set of sine transform coefficients, and subsequently mapping these coefficients to output functions. This universal approximation property is similar to those of neural operators (Kovachki et al., 2023).
@@
Assumption 1: $C_{0}([0,T];{\mathbb{R}}^{p})$ with the $L^{\infty}$-norm is a Banach space. Its metric is defined as ${\lVert{\mathbf{u}_{1}(t)-\mathbf{u}_{2}(t)}\rVert}_{L^{\infty}}$ with $\mathbf{u}_{1}(t)$ and $\mathbf{u}_{2}(t)$ $\in$ $C_{0}([0,T];\mathbb{R}^{p})$.
Assumption 2: K is a compact subset of $C_{0}([0,T];\mathbb{R}^{p})$} and it is closed. By virtue of the Arzelà-Ascoli theorem (Beattie and Butzmann, 2013), all functions belonging to K are uniformly bounded. There exists a constant $B_{K}$ such that all elements of $K$ are bounded by $B_{K}$, that is ${\lVert{u_{i}(t)}\rVert}_{L^{\infty}}\leq B_{K}$ for all $\mathbf{u}(t)\in K$ and $1\leq i\leq p$, where $u_{i}(t)$ is the $i^{\mathrm{th}}$ element of $\mathbf{u}(t)$. It follows from the Arzelà-Ascoli theorem that the functions belonging to $K$ are equicontinuous, and there exists a continuous modulus of continuity $\mathbf{\upphi}_{K}(t)=[\phi_{K,1}(t),\phi_{K,2}(t),...,\phi_{K,p}(t)]^{\top}$ such that for arbitrary $0\leq t_{1}\leq t_{2}\leq T$ and all $\mathbf{u}(t)\in K$, it is satisfied that $|u_{i}(t_{1})-u_{i}(t_{2})|\leq\phi_{K,i}(t_{2}-t_{1})$, thereby $|\mathbf{u}(t_{1})-\mathbf{u}(t_{2})|\leq\left|\mathbf{\upphi}_{K}(t_{2}-t_{1})\right|$, where $\phi_{K,i}(t):[0,+\infty)\to[0,+\infty)$ is monotonic with $\phi_{K,i}(t)\to 0$ as $t\to 0$.